OpenStudy (anonymous):

Solve the inequality \[x^4+5x^2-36 ge 0\] and express the solutions in terms of an interval

OpenStudy (anonymous):

\[x^4-5x^2-36 \ge 0\]

OpenStudy (kinggeorge):

Let \(u=x^2\). Can you solve the quadratic \(u^2-5u-36=0\)?

OpenStudy (anonymous):

yes I figured it out thank you for your response (x^2-4)(x^2+9)

OpenStudy (kinggeorge):

Did you also get where the expression is \(\ge 0\)?

OpenStudy (anonymous):

Solution: \[(-\infty,-2] U [2,\infty)\]

OpenStudy (anonymous):

Did you get the same intervals?

OpenStudy (kinggeorge):

I would check your factorization one more time. I'm getting different intervals, and a very slightly different factorization.

OpenStudy (anonymous):

\[(x-2)(x+2)(x^2+9) \ge 0\]

OpenStudy (anonymous):

so my zero's are 2 and -2... okay, i'll check

OpenStudy (kinggeorge):

Is your original equation\[x^4-5x^2-36 \ge 0\]or\[x^4+5x^2-36 \ge 0\]

OpenStudy (anonymous):

the latter

OpenStudy (kinggeorge):

In that case, I am getting the same intervals. :P

OpenStudy (anonymous):

okay :).